Solving systems by graphing and substitution
y = 2x - 1
x + y = 2
Please show me if this problem is independent, dependent or inconsistent. More importantly I'd like to know the best way to solve future problems like this

Answer by Alan3354(69443)

y = 2x - 1
x + y = 2
-------------
Sub for y in the 2nd eqn
x + (2x-1) = 2
3x - 1 = 2
3x = 3
x = 1
--------
y = 2x - 1
y = 1
------
Since there's a unique solution (1,1) they're independent.
-----------
To graph them, plot 2 points each and draw lines thru them.